The Production Department of Hruska Corporation has submitted the following forecast of units to be produced by quarter for the upcoming fiscal year:

1st Quarter 2nd Quarter 3rd Quarter 4th Quarter
Units to be produced 11,400 10,400 12,400 13,400

Each unit requires 0.30 direct labor-hours and direct laborers are paid $12.50 per hour.

In addition, the variable manufacturing overhead rate is $1.50 per direct labor-hour. The fixed manufacturing overhead is $94,000 per quarter. The only noncash element of manufacturing overhead is depreciation, which is $34,000 per quarter.

Required:

1. Calculate the company’s total estimated direct labor cost for each quarter of the upcoming fiscal year and for the year as a whole. Assume that the direct labor workforce is adjusted each quarter to match the number of hours required to produce the estimated number of units produced.

2&3. Calculate the company’s total estimated manufacturing overhead cost and the cash disbursements for manufacturing overhead for each quarter of the upcoming fiscal year and for the year as a whole.
Answer #1

Answer-1)-

HRUSKA CORPORATION
TOTAL DIRECT LABOR COSTS
Particulars First Quarter Second Quarter Third Quarter First Quarter Year
Direct labor costs 42750 39000 46500 50250 178500

Explanation-

HRUSKA CORPORATION
TOTAL DIRECT LABOR COSTS
Particulars First Quarter Second Quarter Third Quarter First Quarter Year
Units to be produced (a) 11400 10400 12400 13400 47600
Each units requires hours (b) 0.30 0.30 0.30 0.30 0.30
Budgeted direct labor hours (c=a*b) 3420 3120 3720 4020 14280
Direct labor rate per hour (d) $12.50 $12.50 $12.50 $12.50 $12.50
Direct labor costs (e=c*d) $ 42750 39000 46500 50250 178500

2)-

HRUSKA CORPORATION
TOTAL DIRECT LABOR COSTS
Particulars First Quarter Second Quarter Third Quarter First Quarter Year
Total manufacturing overhead $ 99130 98680 99580 100030 397420

Explanation-

HRUSKA CORPORATION
Manufacturing Overhead budget
Particulars First Quarter Second Quarter Third Quarter First Quarter Year
Units to be produced (a) 11400 10400 12400 13400 47600
Each units requires hours (b) 0.30 0.30 0.30 0.30 0.30
Budgeted direct labor hours (c=a*b) 3420 3120 3720 4020 14280
Variable manufacturing overhead rate per DLH (d) $1.50 $1.50 $1.50 $1.50 $1.50
Variable manufacturing overhead (e=c*d) $ 5130 4680 5580 6030 21420
Add:- Fixed manufacturing overhead $ 94000 94000 94000 94000 376000
Total manufacturing overhead $ 99130 98680 99580 100030 397420

3)-

HRUSKA CORPORATION
Manufacturing Overhead budget
Particulars First Quarter Second Quarter Third Quarter First Quarter Year
Cash disbursements for manufacturing overhead $ 65130 64680 65580 66030 261420

Explanation-

Manufacturing Overhead budget
Particulars First Quarter Second Quarter Third Quarter First Quarter Year
Units to be produced (a) 11400 10400 12400 13400 47600
Each units requires hours (b) 0.30 0.30 0.30 0.30 0.30
Budgeted direct labor hours (c=a*b) 3420 3120 3720 4020 14280
Variable manufacturing overhead rate per DLH (d) $1.50 $1.50 $1.50 $1.50 $1.50
Variable manufacturing overhead (e=c*d) $ 5130 4680 5580 6030 21420
Add:- Fixed manufacturing overhead $ 94000 94000 94000 94000 376000
Total manufacturing overhead $ 99130 98680 99580 100030 397420
Less:- Depreciation $ 34000 34000 34000 34000 136000
Cash disbursements for manufacturing overhead $ 65130 64680 65580 66030 261420